[Brahms-l] Shift report 20050604 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sat Jun 04 2005 - 15:59:22 EDT
Shifter leader was sjs
Shifters were ps


Two short stores during shift.  STAR and PHENIX report high backgrounds.
Runs with FS at 2.3 deg, MRS at 52.5 deg. 
(# 15179 - 15191)
0.6M trig 2 events, 0.18 trig 3 events
T3 showing strange behavior (see elog)...problem noted in previous shift.  
FV came in to replay downscale unit.   Had lost ds trig 6 earlier in shift.


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Jun 4 15:59:49 2005

This archive was generated by hypermail 2.1.8 : Sat Jun 04 2005 - 16:00:02 EDT